OpenStudy (anonymous):

find the square root of 1.69? show work

OpenStudy (anonymous):

what work? You just need to put the equation on your calculator ^^ sqrt(1.69) = 1.3

OpenStudy (anonymous):

\[\sqrt{1.69}=\sqrt{\frac{169}{100}}=\frac{13}{10}=1.3\]